# Compression fronts from fast radio bursts

### v1

@@ 3 Plasma heating by the wave
(1) Waves with frequency $\tilde{\omega}$ exceeding the gyrofrequency $\tilde{\omega}_{B}=e\tilde{B}_{\rm bg}/mc$ drive regular particle oscillations with
$\tilde{\gamma}=\sqrt{1+a^{2}}\approx a\gg 1\qquad(\tilde{\omega}_{B}<\tilde{\omega}).$ (9)
Thus, the local $\tilde{\gamma}$ in the packet simply tracks the local strength parameter of the wave $a$.
(2) In the other heating regime, $1<\tilde{\omega}_{B}/\tilde{\omega}<a$, the resonant pumping mechanism is activated, and the wave drives stochastic particle acceleration. Then, the average $\tilde{\gamma}$ inside a fluid element quickly exceeds $a$ and keeps growing as the fluid moves through the wave packet (until radiative losses limit the growth or fluid exits behind the packet). The mean expectation for the particle Lorentz factor $\tilde{\gamma}$ grows according to the equation (B22,Beloborodov 2024)
$\frac{d\tilde{\gamma}}{d\tilde{t}}=\chi\frac{a_{0}^{2}\tilde{\omega}_{B}^{1/3}\tilde{\omega}^{2/3}}{\tilde{\gamma}^{4/3}}\qquad(1<\tilde{\omega}_{B}/\tilde{\omega}<a_{0}),$ (10)
where $a_{0}^{2}=2a^{2}$, $\chi\sim 1$ is a numerical coefficient, and all quantities are measured in the fluid rest frame $\tilde{\cal K}$.
@@
In general, $\overline{\boldsymbol{E}}\neq 0$ which corresponds to a moving MHD fluid $v_{\rm D}\neq 0$. If $v_{\rm D}$ varies (the plasma is accelerated in the wave packet), no global inertial frame exists where the fluid would be static, and one has to consider the heating effect of the wave in a moving MHD fluid. Then, the test particle motion should be examined with $E_{\rm bg}\neq 0$.
Remarkably, the results found in B22 at $E_{\rm bg}=0$ also hold in the accelerating background, as long as the particle motion is measured in the local drift frame $\tilde{\cal K}$ where $\tilde{\boldsymbol{E}}_{\rm bg}=0$. This fact is illustrated below by direct calculations of particle motion in a wave packet with a non-static background field. We will prescribe smoothly varying fields $\boldsymbol{B}_{\rm bg}(\xi)$ and $\boldsymbol{E}_{\rm bg}(\xi)$ (varying on a scale much greater than the particle Larmor radius) and repeat the calculation of B22 for test particles.
In the simulations, we will use the frame where the upstream plasma is at rest, $\beta_{\rm D}=\beta_{\rm u}=0$, and prescribe the background field of the form
$B_{\rm bg}=\frac{\kappa^{2}+1}{2}B_{\rm bg}^{\rm u},\qquad E_{\rm bg}=\frac{\kappa^{2}-1}{2}B_{\rm bg}^{\rm u},$ (12)
The upstream has $\kappa=\kappa_{\rm u}=1$, $E_{\rm bg}=0$, and $B_{\rm bg}=B_{\rm bg}^{\rm u}$. Fluid motion inside the wave packet is specified by a chosen function $\kappa(\xi)$. For the present purposes, Equation (12) may be viewed as an arbitrary prescription to test heating of fluid with a non-uniform motion $\beta_{\rm D}(\xi)$. As explained in next sections, any steady pattern of fluid motion in the wave packet satisfies Equation (12). It also satisfies $\tilde{\rho}/\kappa=const$ (uniform across the packet), where $\tilde{\rho}=\rho/\gamma_{\rm D}$ is the fluid proper density. Thus, the adopted prescription corresponds to a steady flow pattern with the proper compression factor $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}=\kappa$. The magnetic field $\tilde{B}_{\rm bg}$ is frozen in the fluid and compressed by the same factor $\kappa$ while $B_{\rm bg}=\gamma_{\rm D}\tilde{B}_{\rm bg}$ obeys Equation (12).
Our simulation will track a test particle initially located in the static upstream region and exposed to the propagating wave packet. The particle motion in the wave packet is found by solving the dynamical equation,
@@
where vector $\boldsymbol{u}=\gamma\boldsymbol{\beta}$ is the spatial part of the particle four-velocity; radiative losses are neglected here. The fields $\boldsymbol{E}(\xi)$ and $\boldsymbol{B}(\xi)$ include the background and wave fields: $\boldsymbol{E}=\boldsymbol{E}_{\rm bg}+\boldsymbol{E}_{\rm wave}$ and $\boldsymbol{B}=\boldsymbol{B}_{\rm bg}+\boldsymbol{B}_{\rm wave}$. [^4] Notation here differs from B22 where $\boldsymbol{E}$ and $\boldsymbol{B}$ denoted the wave fields and the dynamical equation assumed $\boldsymbol{E}_{\rm bg}=0$.
3.1 Regular heating ($\tilde{\omega}_{B}<\tilde{\omega}$)
@@
$\kappa^{2}=1+\zeta a^{2},$ (15)
@@
where $\zeta$ is a constant coefficient. The prescribed $\kappa(\xi)$ implies the fastest drift of the fluid in the middle of the packet where $a(\xi)$ peaks (Equation 5); the fluid is static ahead and behind the wave packet where $a=0$. The case of $\zeta=0$ corresponds to a static magnetic background throughout the packet ($E_{\rm bg}=0$); in this case, the particle motion is similar to that found in B22. When $\zeta\neq 0$, the presence of $E_{\rm bg}\neq 0$ changes the particle motion by introducing an additional electric force.
After solving for the oscillating particle motion in the wave packet, we calculate its Lorentz factor averaged over the wave oscillation period. Figure 1 shows the obtained Lorentz factor as a function of position $\xi$ in the packet, in three sample models. The models have the same radio wave (packet with amplitude $a_{\max}=4$) and different $\zeta=0$, 0.4, and 1. The results demonstrate that the average particle Lorentz factor $\tilde{\gamma}(\xi)$ measured in the local fluid frame is the same in all three models, independent of the fluid motion in the packet, and follows the solution found with a static background (Equation 9). In the presented numerical models, we chose $\omega_{B}/\omega=0.05$, where $\omega_{B}\equiv\tilde{\omega}_{B}^{\rm u}$ is the gyro-frequency in the upstream plasma. Then, the condition $\tilde{\omega}_{B}<\tilde{\omega}$ is satisfied throughout the packet, so the heating everywhere acts in the regular regime according to Equation (9).
@@
Figure 1: Lorentz factor $\gamma$ of a test particle moving through the wave packet (averaged over the wave oscillation period). The upstream plasma is at rest ($\gamma=1$) and has the gyro-frequency $\omega_{B}=0.05\omega$. The plasma particle moves through the wave packet in the coordinate $\xi=t-z/c$ from right to left. The simulated packet consists of $10^{3}$ oscillations with an envelope described by Equation (5) with $a_{\max}=4$, and the background field is described by Equations (12) and (15). Different colors correspond to models with $\zeta=0$ (static background, $E_{\rm bg}=0$; black), $\zeta=0.4$ (blue), and $\zeta=1$ (green). For each model, dashed curve shows $\gamma$ measured in the fixed frame (where the upstream plasma is at rest), and solid curve shows $\tilde{\gamma}$ measured in the local fluid rest frame. All solid curves overlap, demonstrating that $\tilde{\gamma}(\xi)$ is independent of the local fluid velocity $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$ and follows the solution $\tilde{\gamma}^{2}=1+a^{2}(\xi)$ (red dotted curve).
The test-particle motion changes if $\omega_{B}/\omega$ is increased so that $\tilde{\omega}_{B}/\tilde{\omega}$ exceeds unity somewhere inside the packet, triggering stochastic acceleration. Compression of $\tilde{B}_{\rm bg}$ by the factor of $\kappa$ implies $\tilde{\omega}_{B}=\kappa\omega_{B}$, and the Doppler-shifted wave frequency in the fluid frame is $\tilde{\omega}=\omega/\kappa$. This gives
$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}=\kappa^{2}\,\frac{\omega_{B}}{\omega}.$ (16)
@@
3.2 Stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$)
First, let us formulate the analytical prediction for stochastic heating by solving Equation (10) for $\tilde{\gamma}(\xi)$. The fluid proper time $\tilde{t}$ is related to time $t$ in the lab frame by $d\tilde{t}=dt/\gamma_{\rm D}$. During time $dt$, the fluid element moves in $\xi$ by $d\xi=(1-\beta_{\rm D})dt$, so
@@ 3.2 Stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$)
Using also $\tilde{\omega}=\omega/\kappa$ and $\tilde{\omega}_{B}=\kappa\omega_{B}$ (where $\omega_{B}\equiv\tilde{\omega}_{B}^{\rm u}$ is the upstream gyrofrequency), we rewrite Equation (10) as
$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}=\chi a_{0}^{2}\kappa^{2/3}\left(\omega_{B}\omega^{2}\right)^{1/3}.$ (18)
For any chosen $\kappa(\xi)$, Equation (18) can be integrated for $\tilde{\gamma}(\xi)$, and this analytical prediction can be compared with simulations of test particle motion. Unlike regular heating (where $\tilde{\gamma}$ tracks the rise and drop of $a$), stochastic heating gives a monotonically growing mean expectation $\tilde{\gamma}(\xi)$: $d\tilde{\gamma}/d\xi\geq 0$.
Our simulation setup is described above; in particular, $B_{\rm bg}$ and $E_{\rm bg}$ are given by Equation (12). We arbitrarily choose $\kappa(\xi)$ of the form, [^5] Plasma in packets with monotonically growing $\tilde{\gamma}(\xi)$ is expected to have a monotonically growing $\kappa(\xi)$. Therefore, in sample models of stochastic heating we choose $\kappa(\xi)$ that satisfy $d\kappa/d\xi\geq 0$.
$\kappa=\left\{\begin{array}[]{cc}1+(\kappa_{0}-1)\sin^{2}(\pi\xi/T)&\qquad\xi<T/2\\ \kappa_{0}&\qquad\xi>T/2\end{array}\right.$ (21)
so $\kappa(\xi)$ grows until it reaches $\kappa_{0}$ in the middle of the wave packet and stays constant in the second half of the packet. We have calculated particle motion in packets with $\kappa_{0}=1$ (no fluid motion, $\beta_{\rm D}=0$) and $\kappa_{0}=3$.
@@
Figure 2: Average Lorentz factor $\tilde{\gamma}$ (measured in the local fluid rest frame) of particles crossing the wave packet in the regime of stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$). The average $\tilde{\gamma}$ is calculated by tracking $4\times 10^{3}$ test particles. The upstream plasma is at rest ($\gamma_{\rm D}=1$) and has a small temperature $kT=10^{-2}mc^{2}$; its gyro-frequency is $\omega_{B}=2\omega$. The packet $0<\xi<T$ consists of $N=4\times 10^{3}$ oscillations; its envelope $a_{0}=\sqrt{2}a$ is described by Equation (5) with $a_{0}^{\max}=30$ (blue curve). Background field is described by Equations (12) and (21). As one can see in the figure, the plasma experiences heating as it moves through the wave ($\tilde{\gamma}$ grows with $\xi$). Black solid curves show the simulation results for $\kappa_{0}=1$ (static background field, no bulk motion of the plasma) and $\kappa_{0}=3$ ($\beta_{\rm D}^{\max}=0.8$). The results approximately agree with the analytical prediction (Equation (18), red dotted curves) with $\chi=0.66$ for $\kappa_{0}=1$ and $\chi=0.67$ for $\kappa_{0}=3$.
The stochastic nature of heating requires calculations of many realizations of particle motion in the packet. We draw 4000 particles from an initial Maxwellian distribution with a small temperature $kT=10^{-2}mc^{2}$, seeding small initial perturbations for the stochastic behavior. The particle motion is followed for each of the 4000 realizations, and the result is averaged to find the mean expectation for $\tilde{\gamma}(\xi)$.
The results are shown in Figure 2 for $\omega_{B}/\omega=2$ and $a_{0}=30$. In both shown cases, $\kappa_{0}=1$ and $\kappa_{0}=3$, the found $\tilde{\gamma}(\xi)$ is in good agreement with the analytical prediction given by Equation (18) with $\chi\approx 2/3$.
In summary, the results of B22 for regular and stochastic heating hold in the local frame $\tilde{\cal K}$ regardless of fluid motion, as long as the fluid speed varies on timescales much longer than the Larmor period of particles. The Lorentz factor $\tilde{\gamma}\gg 1$ implies a strong heating effect, giving the fluid a relativistic effective temperature. The energized particles gyrate in the plane perpendicular to $\boldsymbol{B}_{\rm bg}$, so the plasma behaves as a two-dimensional relativistic gas. Its internal energy density $U_{\rm p}$ and pressure $P_{\rm p}$ (measured in frame $\tilde{\cal K}$) are
$U_{\rm p}=2P_{\rm p}=(\tilde{\gamma}-1)\tilde{\rho}c^{2},\qquad\tilde{\rho}=\frac{\rho}{\gamma_{\rm D}}.$ (22)
The heated particles gyrate with the Larmor frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}\ll\tilde{\omega}_{B}$. It is below the wave frequency $\tilde{\omega}$ in both heating regimes, $\tilde{\omega}_{B}<\tilde{\omega}$ and $\tilde{\omega}_{B}>\tilde{\omega}$ (B22).
4 MHD equations for fluid motion

### v2

@@ III Plasma heating by the radio wave
$\frac{d\langle\tilde{\gamma}\rangle_{\rm ens}}{d\tilde{t}}=\chi\frac{a^{2}\tilde{\omega}_{B}^{1/3}\tilde{\omega}^{2/3}}{\langle\tilde{\gamma}\rangle_{\rm ens}^{4/3}},$ (15)
[…]
@@
In general, $\overline{\boldsymbol{E}}\neq 0$ which corresponds to a moving MHD fluid $v_{\rm D}\neq 0$. If $v_{\rm D}$ varies (the plasma is accelerated in the wave packet), no global inertial frame exists where the fluid would be static, and one has to consider the heating effect of the radio wave in a moving MHD fluid. Then, the test particle motion should be examined with $E_{\rm bg}\neq 0$, as done in this section.
One goal of our calculations below is to generalize the threshold condition for stochastic heating $b\equiv\tilde{\omega}_{B}/\tilde{\omega}\gtrsim 1$ so that it applies to accelerating fluids. If the upstream plasma has $b_{\rm u}\gtrsim 1$, stochastic heating is triggered immediately in the leading part of the wave packet once $a$ exceeds $\sim 1$ (B22). If the upstream plasma has $b_{\rm u}\ll 1$, stochastic heating can still occur, but deeper inside the wave packet where $b$ is sufficiently increased by fluid acceleration and compression. This can occur after the wave strength parameter has reached a high $a\gg 1$, which changes the threshold for stochastic heating.
Remarkably, in an accelerating fluid, Equations (14) and (15) still hold, i.e. the evolution of $\tilde{\gamma}$ in both regimes remains unchanged. This fact will be illustrated below by direct calculations of particle motion in a wave packet with a non-static background field, which defines the fluid motion. We will prescribe smoothly varying fields $\boldsymbol{B}_{\rm bg}(\xi)$ and $\boldsymbol{E}_{\rm bg}(\xi)$ (varying on a scale much greater than the particle Larmor radius) and repeat the calculation of B22 for test particles.
The background fields will be parameterized as follows
$B_{\rm bg}=\frac{\kappa^{2}+1}{\kappa_{\rm u}^{2}+1}B_{\rm bg}^{\rm u},\qquad E_{\rm bg}=\frac{\kappa^{2}-1}{\kappa_{\rm u}^{2}+1}B_{\rm bg}^{\rm u}.$ (17)
In the simulations, we will use the frame where the upstream plasma is at rest; then, $E_{\rm bg}^{\rm u}=0$ and $\kappa_{\rm u}=1$. Fluid acceleration inside the wave packet is specified by a chosen function $\kappa(\xi)$. For the present purposes, Equation (17) may be viewed as an arbitrary prescription to test heating of fluid with non-uniform motion $\beta_{\rm D}(\xi)$, which is related to $\kappa(\xi)$ by Equation (10). As explained in next sections, any steady pattern of fluid motion in the wave packet actually satisfies Equation (17). It also satisfies $\tilde{\rho}/\kappa=const$ (uniform across the packet), where $\tilde{\rho}=\rho/\gamma_{\rm D}$ is the fluid proper density. Thus, the adopted prescription corresponds to a steady flow pattern with the proper compression factor $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}$. The magnetic field $\tilde{B}_{\rm bg}$ is frozen in the fluid and compressed by the same factor $\kappa/\kappa_{\rm u}$ while $B_{\rm bg}=\gamma_{\rm D}\tilde{B}_{\rm bg}$ obeys Equation (17).
Our simulations will track test particles initially located in the static, cold upstream region exposed to the propagating wave packet. The particle motion in the wave packet is found by solving the dynamical equation,
@@
where vector $\boldsymbol{u}=\gamma\boldsymbol{\beta}$ is the three-dimensional spatial part of the particle four-velocity. Radiative losses are neglected in Equation (18). The fields $\boldsymbol{E}(\xi)$ and $\boldsymbol{B}(\xi)$ include the background and wave fields: $\boldsymbol{E}=\boldsymbol{E}_{\rm bg}+\boldsymbol{E}_{\rm wave}$ and $\boldsymbol{B}=\boldsymbol{B}_{\rm bg}+\boldsymbol{B}_{\rm wave}$. [^5] Notation here differs from B22 where $\boldsymbol{E}$ and $\boldsymbol{B}$ denoted the wave fields and the dynamical equation assumed $\boldsymbol{E}_{\rm bg}=0$.
III.1 Regular oscillations
@@
$\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}=1+\zeta a^{2},$ (19)
@@
where $\zeta$ is a constant coefficient. The simulation will be done in the reference frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$). The case of $\zeta=0$ corresponds to a static magnetic background throughout the packet ($\kappa=1$, $E_{\rm bg}=0$); in this case, the particle motion is similar to that found in B22. When $\zeta\neq 0$, the increasing $\kappa>1$ means $\beta_{\rm D}>0$ and the appearance of additional electric force $eE_{\rm bg}$, which changes the particle motion. The prescribed $\kappa(\xi)$ gives the fastest drift of the fluid in the middle of the packet where $a(\xi)$ peaks; the fluid is static ahead and behind the wave packet where $a=0$.
We have calculated three sample models with $\zeta=0$, 0.4, and 1. In all three models, the radio wave packet has the amplitude $a_{\max}=4$, and the upstream plasma has the gyro-frequency $\tilde{\omega}_{B}^{\rm u}=0.03\omega$, where $\omega$ is the wave frequency measured in the upstream rest frame. These parameters represent an FRB propagating through the magnetar wind at radii $r\sim 3\times 10^{12}$ cm, as discussed in section VII below. Another parameter is the number of wave oscillations in the packet, $N=\omega T/2\pi$. Real FRBs have $N\sim 10^{6},$ and the presented simulations have $N=10^{3}$. We have also run models with different $N$ and verified that the results are independent of $N$ as long as $N\gg a\tilde{\omega}/\tilde{\omega}_{B}$ (this is simply the condition that the packet is much wider than the particle Larmor radius).
For each model, we have solved for the particle motion in the radio wave packet, and then calculated its Lorentz factor averaged over the wave oscillation period: $\gamma$ in the lab frame and $\tilde{\gamma}$ in the local fluid rest frame. Figure 1 shows the average $\gamma$ and $\tilde{\gamma}$ as functions of position $\xi$ in the packet. It demonstrates that $\tilde{\gamma}=\sqrt{1+a^{2}}$ holds in all three models ($\zeta=0$, 0.4, 1). In the fluid frame, the particle performs same regular oscillations, independent of the fluid bulk acceleration in the packet. Since fluid motion $\kappa$ varies on scales much larger than the Larmor scale, Larmor rotation is not excited (B22). The particle orbit measured in the local fluid frame is nearly identical to the canonical “8”-shaped orbit, same as found in the absence of any background field (Appendix B).
@@
Figure 1: Lorentz factor $\gamma$ of a test particle moving through the wave packet (averaged over the wave oscillation period). The upstream plasma is cold ($\tilde{\gamma}=1$) and static ($\kappa_{\rm u}=1$); it has gyro-frequency $\tilde{\omega}_{B}^{\rm u}=0.03\omega$ measured in the fluid frame. The test particle moves through the wave packet in the coordinate $\xi=t-z/c$ from right to left. The simulated packet consists of $10^{3}$ oscillations with an envelope described by Equation (7) with $a_{\max}=4$, and the background field is described by Equations (17) and (19). Different colors correspond to models with $\zeta=0$ (static background, $E_{\rm bg}=0$; black), $\zeta=0.4$ (blue), and $\zeta=1$ (green). For each model, dashed curve shows $\gamma$ measured in the fixed frame (where the upstream plasma is at rest), and solid curve shows $\tilde{\gamma}$ measured in the local fluid rest frame. All three solid curves overlap, demonstrating that $\tilde{\gamma}(\xi)$ is independent of the local fluid velocity $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$ and follows the solution $\tilde{\gamma}^{2}=1+a^{2}(\xi)$ (red dotted curve, overlaping the solid curves).
III.2 Stochastic heating
The onset of stochastic heating is controlled by two parameters: $a$ and
$b\equiv\frac{\tilde{\omega}_{B}}{\tilde{\omega}}$. (20)
As an example, consider a radio wave packet with amplitude $a_{\max}=30$ propagating in an ambient plasma with gyrofrequency $\tilde{\omega}_{B}^{\rm u}=0.2\tilde{\omega}_{\rm u}$, so the upstream plasma has $b_{\rm u}=0.2$. These parameters may represent an FRB propagating through a magnetar wind at $r\sim 5\times 10^{11}$ cm (section VII). In this case, the fluid is accelerated and compressed inside the wave packet by a large factor $\kappa/\kappa_{\rm u}\gg 1$, and $b\gg 1$ can be reached inside the packet. Indeed, Equation (17) implies compression of $\tilde{B}_{\rm bg}$ by the factor of $\kappa/\kappa_{\rm u}$, so the gyrofrequency measured in the local fluid frame $\tilde{\cal K}$ is increased to $\tilde{\omega}_{B}=(\kappa/\kappa_{\rm u})\tilde{\omega}_{B}^{\rm u}$. At the same time, the wave frequency measured in the fluid frame is reduced from $\tilde{\omega}_{\rm u}=\omega/\kappa_{\rm u}$ to $\tilde{\omega}=\omega/\kappa$. This gives $b=(\kappa/\kappa_{\rm u})^{2}b_{\rm u}$.
The threshold condition for the onset of stochastic heating can be stated in the form $b>b_{\rm s}(a)$, as discussed in section III.3 below. The transition to stochastic heating occurs for any profile of $\kappa(\xi)$ as long as it reaches $b>b_{\rm s}$, i.e. $\kappa/\kappa_{\rm u}$ exceeds $(b_{\rm s}/b_{\rm u})^{1/2}$. In our example simulations, we prescribe $\kappa(\xi)$ using a modified version of Equation (19): [^6] We use the simplest modification of Equation (19) that satisfies $d\kappa/d\xi\geq 0$. This condition is motivated by the results of later sections that describe self-consistent calculations of $\kappa(\xi)$. Unlike regular oscillations (where $\tilde{\gamma}$ tracks the rise and drop of $a(\xi)$ in the wave packet), stochastic heating gives a monotonically growing expectation value $\tilde{\gamma}(\xi)$: $d\tilde{\gamma}/d\xi\geq 0$, leading to $d\kappa/d\xi\geq 0$.
$\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}=\left\{\begin{array}[]{cc}\vskip 2.84526pt1+\zeta a^{2}&\qquad\xi<T/2\\ 1+\zeta a_{\max}^{2}&\qquad\xi>T/2\end{array}\right$. (23)
The results will be shown for two sample models: $\zeta=0$ (no fluid acceleration in the packet) and $\zeta=1$.
Measurements of stochastic heating require calculations of many realizations of particle motion in the wave packet. We draw 4000 upstream particles from an initial Maxwellian distribution with a small temperature $kT=10^{-2}mc^{2}$, seeding initial perturbations for the stochastic behavior. The particle motion is followed for each of the 4000 realizations, and the result is averaged to find the mean value $\tilde{\gamma}(\xi)$. The results are shown in Figure 2.
In the model with $\zeta=0$, the ratio $b\equiv\tilde{\omega}_{B}/\tilde{\omega}=0.2$ stays uniform and small across the wave packet. Then, the stochastic heating does not occur: the plasma everywhere performs regular oscillations with $\tilde{\gamma}=\sqrt{1+a^{2}}$. By contrast, in the model with $\zeta=1$, the background field is subject to strong compression, $\kappa/\kappa_{\rm u}=\sqrt{1+a^{2}}$, which changes the particle response to the radio wave. In the leading part of the packet, where $\kappa/\kappa_{\rm u}$ is still small, the particles perform regular oscillations with $\tilde{\gamma}=\sqrt{1+a^{2}}$, same as in the model with $\zeta=0$. The transition to stochastic heating is observed at $\xi_{\rm stoch}\approx 140\times 2\pi/\omega$, where $\tilde{\gamma}(\xi)$ shows a steep growth.
@@ III.2 Stochastic heating
Using also $\tilde{\omega}=\omega/\kappa$ and $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$, we rewrite Equation (15) as
$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}=\chi\,a^{2}\kappa^{2/3}\left(\frac{\tilde{\omega}_{B}^{\rm u}\omega^{2}}{\kappa_{\rm u}}\right)^{1/3}.$ (25)
For any chosen $\kappa(\xi)$, Equation (25) can be integrated for $\tilde{\gamma}(\xi)$. As one can see in Figure 2, this analytical calculation reproduces the mean value $\tilde{\gamma}(\xi)$ found in the full dynamical simulation with the ensemble of 4000 particles. We have also calculated models with different $a_{\max}$, $\tilde{\omega}_{B}^{\rm u}/\tilde{\omega}_{\rm u}$, and $\zeta$, and found that Equation (25) with $\chi\approx 0.8$ well describes stochastic heating in all of them.
@@
Figure 2: Average Lorentz factor $\tilde{\gamma}$ (measured in the local fluid rest frame) of particles interacting with the wave packet with amplitude $a_{\max}=30$. The upstream plasma is at rest ($\kappa_{\rm u}=1$) and has a small temperature $kT=10^{-2}mc^{2}$; its gyro-frequency is $\tilde{\omega}_{B}^{\rm u}=0.2\tilde{\omega}_{\rm u}$. The average $\tilde{\gamma}$ is calculated by tracking $4\times 10^{3}$ test particles. Background fields $B_{\rm bg}$ and $E_{\rm bg}$ are described by Equations (17) and (23). The model with $\zeta=0$ corresponds to no fluid acceleration in the wave packet and shows regular oscillations with $\tilde{\gamma}=\sqrt{1+a^{2}}$. The model with $\zeta=1$ demonstrates the transition to stochastic heating caused by the strong acceleration and compression of the fluid. Red dotted curves show the analytical predictions for regular oscillations ($\tilde{\gamma}=\sqrt{1+a^{2}}$) and stochastic heating (Equation (25) with $\chi=0.8$).
Figure 3: Transition curve separating the regimes of regular oscillations and stochastic heating on the $a$-$b$ plane. Black circles show the results of 12 simulations: each circle shows $(a,b)$ at which the transition from regular oscillations to stochastic heating occurred as the particles moved through the wave packet. Dotted curve shows $b=(1/3)\sqrt{1+a^{2}}$. The region above the solid curve satisfies $b>2\sqrt{2}\,a$, which corresponds to wave amplitude $\tilde{E}_{0}<\tilde{B}_{\rm bg}/2$. In this region, the electromagnetic field satisfies $E^{2}<B^{2}$ throughout the radio wave oscillation, and stochastic heating is not expected. The radio wave cannot compress the background far above the solid line, as this would require energy exceeding the available energy of the wave.
As illustrated in Figures 1 and 2, the particle response to the radio wave admits a simple analytical description in frame $\tilde{\cal K}$ regardless of how this frame moves in the radio wave packet. This universality is foreseeable. The speed of frame $\tilde{\cal K}$ varies on timescales much longer than the particle Larmor period; so, in the problem of particle response to the radio wave, $\tilde{\cal K}$ is not too different from an inertial frame. As a result, $\tilde{\gamma}$ evolves with the same rate as in a non-accelerating fluid: it follows Equation (14) in the regular oscillation regime and Equation (25) in the stochastic heating regime.
The large Lorentz factor $\tilde{\gamma}\gg 1$ may be thought of as a heating effect, giving the fluid a relativistic effective enthalpy and increasing its effective inertial mass. Note that the Larmor frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}$ remains below the wave frequency $\tilde{\omega}$ for both regular oscillation and stochastic heating regimes (B22). The ratio $\tilde{\omega}_{\rm L}/\tilde{\omega}$ reaches a maximum value of $\approx 1/3$ at the transition between the two regimes, which is described below.
III.3 Transition curve on the $a$-$b$ plane
Each test-particle simulation that shows the transition to stochastic heating inside the wave packet at some point $\xi_{\rm stoch}$ gives a measurement of $a(\xi_{\rm stoch})$ and $b(\xi_{\rm stoch})$. These measurements are shown in Figure 3 for a set of simulations with various $b_{\rm u}$, $a_{\max}$, and $\zeta$. We kept $b_{\rm u}$ below 0.4 in all cases; then, stochastic heating is not activated without accelerating/compressing the background. Compression enables the transition by increasing $b$ from $b_{\rm u}$ by the factor $b/b_{\rm u}=\kappa^{2}/\kappa_{\rm u}^{2}=1+\zeta a^{2}$.
As one can see in Figure 3, the transition occurs on a well defined curve in the $a$-$b$ plane. It is independent of the shape of the wave packet $a(\xi)$ or the profile of background compression in the packet (here parameterized by $\zeta$); it is also independent of $b_{\rm u}$. Basically, the transition from regular oscillations to stochastic heating is controlled by local $a$ and $b$. This is expected, since the local plasma response to the wave viewed in the fluid frame is independent of the drift of this frame or the global shape of the wave packet.
The transition curve extends from $a\sim 0.3$ to $a\gg 1$, and approximately follows the relation
$b=b_{\rm s}\approx\frac{1}{3}\sqrt{1+a^{2}}\qquad{\rm(transition)}.$ (26)
At $b<b_{\rm s}$, the plasma performs regular oscillations in the radio wave with frequency $\tilde{\omega}$ while Larmor rotation (with frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}=\tilde{\omega}_{B}/\sqrt{1+a^{2}}<\tilde{\omega}/3$) is not excited. Excitation of Larmor rotation and transition to stochastic heating occur when $\tilde{\omega}_{\rm L}$ exceeds $\tilde{\omega}/3$ (Larmor radius $\tilde{r}_{\rm L}=c/\tilde{\omega}_{\rm L}$ exceeds half wavelength $\pi c/\tilde{\omega}$), which corresponds to $b>b_{\rm s}$. At this point, the quick growth of $\tilde{\gamma}$ is triggered, with $\tilde{\gamma}$ reaching very high values and $\tilde{\omega}_{\rm L}$ decreasing far below $\tilde{\omega}$.
Within a numerical factor, Equation (26) can be obtained by combining $\tilde{\gamma}\sim a$ (achieved in regular oscillations) with the condition for triggering stochastic heating $\tilde{\gamma}\gtrsim\tilde{\gamma}_{\star}\approx(a^{3}/b)^{1/2}$ (B22). The transition to stochastic heating at $\tilde{\gamma}\sim\tilde{\gamma}_{\star}$ may be thought of as the onset of synchrotron absorption of the wave.
IV MHD equations for fluid motion